# Learning to Construct Implicit Communication Channel

### v1

@@ Abstract
Effective communication is an essential component in collaborative multi-agent systems. Situations where explicit messaging is not feasible have been common in human society throughout history, which motivate the study of implicit communication. Previous works on learning implicit communication mostly rely on theory of mind (ToM), where agents infer the mental states and intentions of others by interpreting their actions. However, ToM-based methods become less effective in making accurate inferences in complex tasks. In this work, we propose the Implicit Channel Protocol (ICP) framework, which allows agents to construct implicit communication channels similar to the explicit ones. ICP leverages a subset of actions, denoted as the scouting actions, and a mapping between information and these scouting actions that encodes and decodes the messages. We propose training algorithms for agents to message and act, including learning with a randomly initialized information map and with a delayed information map. The efficacy of ICP has been tested on the tasks of Guessing Number, Revealing Goals, and Hanabi, where ICP significantly outperforms baseline methods through more efficient information transmission.
@@ 1 Introduction
Explicit communication uses direct channels independent of the environment dynamics (Sukhbaatar et al., 2016; Foerster et al., 2016; Jiang & Lu, 2018), allowing agents to transmit observations, intentions, and advice to facilitate decision-making and coordination (Zhu et al., 2022; Qu et al., 2021). This approach, analogous to human language or verbal exchanges (Havrylov & Titov, 2017; Baker et al., 1999), has been widely employed in MARL to enhance collaboration. However, dependence on direct channels introduces significant computational and memory overheads (Roth et al., 2006), which makes it challenging to implement in certain scenarios, like tasks without communication channels or decentralized frameworks (Oliehoek et al., 2008; Kraemer & Banerjee, 2016).
@@
To address these challenges associated with ToM methods, we introduce Implicit Channel Protocol (ICP), a novel framework that allows agents to construct communication protocols in implicit communication likewise how it was done in explicit communication. ICP leverages a subset of actions, denoted as the scouting actions, which have no or uniform effects on environment dynamics. A centralized mapping between information and these scouting actions is established to encode and decode the messages. Agents exchange information by deliberately taking scouting actions, forming an implicit communication channel. We further demonstrate how agents’ strategies are trained on this channel, including training with a randomly initialized information map and training with a delayed information map.
We validate the effectiveness of ICP through comprehensive experiments on the tasks of Guessing Number, Revealing Goals, and Hanabi (Bard et al., 2020). These environments share a common characteristic: they lack direct communication but agents must collaboratively make decisions to achieve shared rewards. This setting introduces significant challenges, including sparse and delayed reward feedback, along with difficulty in credit assignment both temporally and among agents. Despite these hurdles, our experiments on Guessing Number and Revealing Goals demonstrate that ICP significantly enhances performance, through more efficient information transmission, compared to baseline methods. In Hanabi, which is a popular card game played by humans, our approach achieved an average score of $24.91$ out of $25$, which surpasses the best available learning algorithm who obtains $23.81$.
@@ 2 Background and Related Work
Cooperative multi-agent problem can be formulated as a Dec-POMDP game (Bernstein et al., 2002), which is described by the tuple $G=(S,U,N,\mathcal{T},O,R,\gamma)$, where $S$ represents the true global state of the environment. At each discrete time step $t$, every agent $i\in N:=\{1,\ldots,n\}$ selects an action $u_{i}$ from its action space $U_{i}$. The state transition function $\mathcal{T}(s^{\prime}|s,u):S\times U\times S\rightarrow P(S)$ governs the transition of states, where $u=(u_{1},\ldots,u_{n})$ denotes the joint action. In POMDP, the global state remains inaccessible, and each agent $i$ can only perceive its individual observation $o_{i}$ through the observation function $O_{i}(s):S\times N\rightarrow O$. $R_{i}(s,u_{i}):S\times U_{i}\rightarrow R$ represents the reward function for each agent $i$. In the cooperative scenarios, all agents share a common reward function $R(s,u):S\times U\rightarrow R$, known as team reward. The objective for each agent is to maximize the expected return, which makes effective cooperation among all agents necessary.
@@
Communication among agents is critical for effective collaboration in MARL. Most researches in this field focused on explicit communication protocols (Tucker et al., 2022; Peng et al., 2017; Kong et al., 2017; Pesce & Montana, 2020; Kim et al., 2019; Wang et al., 2020; Freed et al., 2020b, a; Gupta et al., 2023; Foerster et al., 2016), where agents exchange messages containing critical information. Additionally, techniques like attention mechanisms or graph-based methods that used to build more effective communication connection (Jiang & Lu, 2018; Das et al., 2019; Jiang et al., 2018; Sukhbaatar et al., 2016; Chen et al., 2024; Tucker et al., 2022) and intention sharing or theory of mind (ToM) reasoning to transmit more useful information (Wang et al., 2021; Kim et al., 2020; Qu et al., 2021), further enhancing the effectiveness of MARL systems.
Explicit communication on direct channel often incurs significant communication overhead and may not be available in environments with limited bandwidth or high communication costs. Implicit communication, where agents convey information through their behaviors, has been explored as an alternative approach (Li et al., 2024; Tian et al., 2023; Li et al., 2021; Shaw et al., 2022; Grupen et al., 2022). Theory of mind (ToM)-based methods represent one main research strand for implicit communication, where agents infer the intentions and beliefs of others to anticipate actions and adjust strategies accordingly. Methods in Rabinowitz et al. (2018); Tian et al. (2020); Nguyen et al. (2020); Zhao et al. (2023) aim to enhance coordination by modeling and predicting others’ behavior based on observed actions. However, ToM-based approaches can be computationally intensive and may struggle to accurately infer intentions in complex environments.
@@ 3 Setting
In an idealized situation, scouting actions serve to influence the observation mapping function, which allows the agents to gather critical information from the environment to improve their decision-making and coordination. A scouting action might reveal information about an agent’s surroundings or the state of other agents without changing environment unpredictable. For instance, in the game of Hanabi, Hint actions reveal information to other players while consuming an information token, thereby uniformly altering the game state. Similarly, in StarCraft II, scouting the map with Scan from Terrans requires depleting energy, which impacts the environment and provides new information in the agent’s observations.
@@ 4 Information Carried by Scouting Actions
Scouting actions carry two types of information:
@@
    Information reflected through the environment: When an agent performs a scouting action, it influences the observation function, and the resulting observations provide information. This type of information depends on the observation function, states, and the joint actions of all agents. All scouting action will change observation function and bring information.
@@
    Information reflected through the choice of the scouting action: The specific scouting action chosen by an agent can carry intentional meaning. This type of information only depends on other agents’ scouting action policy. But while receiver agent can’t understand sender agent’s intention, no information will be transmitted.
General multi-agent reinforcement learning (MARL) methods tend to focus more on the first type of information. In Dec-POMDPs without model other agents’ strategies, both partial observe of states and other agents’ policy introduce uncertainty. And in exploration phase of online MARL training, second type of information can’t be utilized due to unknown policies of others. Therefore, MARL methods mostly learn to use the information reflected through the environment.
For second type of information which is not limited by observation function and states, when agents’ intention is understand correctly, it is more fixable and could be more useful for decision making and cooperation. Theory of mind (ToM) methods utilized second type of information by modeling strategies or inferring intention of other agents. By estimating the internal states of others, an agent can interpret the intentional choices behind their scouting actions. One approach is to build an explicit belief model that infers the state from observed actions to achieve communication. However, this method incurs a computational cost that grows exponentially with the size of the game and the number of agents. Moreover, the precise state distribution information is often difficult to fully utilize. Another approach involves training a neural network to map actions to states or intentions, which requires exponentially more data compared to action policy training. Since this method does not produce explicit inferences, the resulting information is inherently biased, and this bias can negatively impact the agent’s decision-making process.
To take advantage of second type information conveyed by scouting actions and to mitigate the computational overhead and inaccuracies of ToM methods, we propose a simpler and more effective solution. By establishing a shared communication protocol where all agents follow the same strategy for selecting scouting actions, the intended information can be easily decoded from the chosen scouting action. This approach eliminates the need for state estimation and intention inference required in ToM methods. In our setting, this communication protocol has following benefits:
@@
    Low Computational Overhead and Accurate Information Transmission: By having all agents follow the same strategy for selecting scouting actions, there is no need for additional computational inference or training of a separate inference model to deduce intentions. The inverse of the strategy can be used to directly interpret the information, resulting in low computational overhead and accurate transmission of intentions.
@@
    Broadcasted Information: Since scouting actions are observable by all agents within the system, any information conveyed through these actions is inherently broadcasted to all participants.
@@
    Independence from Observation-Based Information: When different embeddings are used, the information conveyed by altering the observation function may differ, making the information carried by scouting actions not necessarily correlated with the environment-based information. Both types of information can be optimized concurrently to enhance cooperation and decision-making.
@@ 5 Strategy Training on Constructed Implicit Channel
To establish a shared communication protocol, we propose Implicit Channel Protocol (ICP), a new framework for agents’ implicit communication. In this framework, at each step, an agent chooses whether to send information. If not, a regular action $u^{r}$ is taken according to the action policy. If otherwise the agents wish to convey message $m_{i}\in M$, it maps $m_{i}$ through the centralized mapping mechanism $\mathcal{P}$, which outputs a scouting action $u^{s}=\mathcal{P}(m_{i})$. This mechanism $\mathcal{P}$ needs to be decodable, which provides an output message from scouting actions. By allowing agents to send information through encoded scouting actions and other agents to find the message from the globally observed scouting actions, an implicit communication channel is constructed.
@@
Equation (1) highlights three essential components of the framework: the agents’ action policy $\pi$, the message strategy $\phi$, and the centralized mapping mechanism $\mathcal{P}$. To develop these components, we investigate two different methodologies: The training with random initial information map and the training with delayed information map. Both approaches construct an implicit communication channel within the environment, with which the agents exchange information and achieve implicit coordination. In the rest of the section, we will discuss the details of these approaches.
@@ 5.1 Strategy Training with Random Initial Map
In this approach, we first establish a one-to-one mapping between information and actions using a randomly initialized embedding, that makes the size of the message same as the size of the scouting action space $M=\{1,...,|U^{s}|\}$. Once this one-to-one mapping is set, the inverse of $\mathcal{P}$ will obtain the original message $m_{i}=\mathcal{P}^{-1}(u_{i}^{s})$.
In our implementation, we use a value based method for training the action policy, incorporating parameter sharing and value decomposition techniques to facilitate effective coordination among agents. For the information strategy, we use the communication gradient method, which is designed to optimize communication protocols within a predefined channel. Both the action policy and the information strategy are parameterised as Q-networks, for action policy $\pi_{i}(\tau_{t,i})=\text{argmax }Q^{\theta_{1,i}}(\tau_{t,i},u_{i})$, and message strategy $\phi_{i}(\tau_{t,i})=\text{argmax }Q^{\theta_{2,i}}(\tau_{t,i},m_{i})$. Given the value function $V^{\pi_{i},\phi_{i},\mathcal{P}}(s)$, the Q-functions are defined by $Q^{\pi_{i}}(\tau_{t,i},u_{i}):=\mathbb{E}_{m\sim\phi,\mathcal{P}}[r_{t}(s_{t},u_{t},s_{t+1})+\gamma V^{\pi_{i},\phi_{i},\mathcal{P}}],$ and $Q^{\phi_{i}}(\tau_{t,i},m_{i}):=\mathbb{E}_{u^{r}\sim\pi,\mathcal{P}}[r_{t}(s_{t},u_{t},s_{t+1})+\gamma V^{\pi_{i},\phi_{i},\mathcal{P}}],$ where $\pi=(\pi_{1},...,\pi_{n})$, $\phi=(\phi_{1},...,\phi_{n})$ and $m=(m_{1},...,m_{n})$.
@@
This method ensures a continuous relaxation between $\phi$ and $\vec{m}$. As a result of this relaxation, the game becomes fully differentiable and can be trained using the backpropagation algorithm. By enabling gradients to flow smoothly through the communication process, this approach facilitates more effective and efficient training of the agents. Compare the communication gradient implementation in DIAL (Foerster et al., 2016), we adopt this Gumbel-Softmax technique to replace the Discretize/Regularize Unit (DRU). This substitution allows the agents to send discrete and non-binary messages during the learning phase, thereby enhancing the richness and accuracy of the information exchanged. Consequently, this method simplifies the learning process and promotes better coordination and communication among the agents.
@@ 5.2 Strategy Training with Delayed Map
In the training process with delayed map, we begin by learning the information strategy and action policy using an direct communication channel. This initial phase provides agents with a flexible environment that allows for direct communication, enabling them to develop their strategies more effectively. Once these strategies are well-established, we then define the mapping mechanism that maps the messages to actions. Finally, we fine-tune both the action policies and information strategies in ICP framework to ensure optimal performance within the constraints of the final environment.
Given that the information strategy was initially trained using an direct communication channel, there may be discrepancies between this channel and the implicit communication channel constructed by scouting actions, particularly in terms of capacity and whether the channel is discrete or noiseless. To minimize any potential performance loss resulting from these differences, we can limit the capacity of the direct channel or apply efficient embedding techniques during the message-to-action mapping process. These steps help to align information strategies developed in the direct channel with the constraints of the implicit channel and ensure messages are decodable.
During implementation, we utilize standard explicit communication algorithms designed for direct channels, such as RGMComm, to pre-train the information strategy. Then we use two kinds of mapping methods, including simple one-to-one mapping and the hat mapping method to obtain new information strategies. The one-to-one mapping ensures that each message corresponds directly to an action, while the hat mapping allows an agent to communicate efficiently with multiple receivers. Finally, we fine-tune both the information strategy and the action policy to optimize performance within the implicit communication framework.
@@
In the pre-training phase, we utilize an direct communication channel to facilitate the learning of the communication strategy. During this phase, the message size can be larger than the scouting action space, allowing for more complex and detailed communication between agents. This flexibility enables the agents to explore a wider range of communication strategies and coordination behaviors, which are crucial for solving complex tasks in partially observable environments. However, after pre-training, the message strategy needs to transition to the implicit communication framework and it is necessary to lossless compress these messages. For one-to-one mapping, this mapping requires that message size match the scouting action space exactly. It ensures a direct and consistent mapping between the learned messages and the corresponding actions.
@@
When the local observations of different agents’ have large overlaps, hat principle-based mapping method utilizes broadcast channels to achieve a single transmission of messages to multiple receivers. The method is inspired by the multi-color hat guessing game, where players use logical deduction based on limited communication to maximize team success. In this game, players can observe the hat colors of others but not their own, and they must guess their own hat color using a pre-defined strategy. A well-known strategy involves each player leveraging the sum of the hat colors they observe, modulo the total number of possible colors. For example, in an 8-color version of the game, each player can compute the sum of the colors they observe (assigned numerical values) and use a pre-determined rule, such as calculating the sum modulo 8, to make a logical deduction about their own hat color.

### v2

# Learning to Communicate Through Implicit Communication Channels

@@ Abstract
Effective communication is an essential component in collaborative multi-agent systems. Situations where explicit messaging is not feasible have been common in human society throughout history, which motivate the study of implicit communication. Previous works on learning implicit communication mostly rely on theory of mind (ToM), where agents infer the mental states and intentions of others by interpreting their actions. However, ToM-based methods become less effective in making accurate inferences in complex tasks. In this work, we propose the Implicit Channel Protocol (ICP) framework, which allows agents to communicate through implicit communication channels similar to the explicit ones. ICP leverages a subset of actions, denoted as the scouting actions, and a mapping between information and these scouting actions that encodes and decodes the messages. We propose training algorithms for agents to message and act, including learning with a randomly initialized information map and with a delayed information map. The efficacy of ICP has been tested on the tasks of Guessing Numbers, Revealing Goals, and Hanabi, where ICP significantly outperforms baseline methods through more efficient information transmission.
@@ 1 Introduction
Explicit communication uses direct channels independent of the environment dynamics (Sukhbaatar et al., 2016; Foerster et al., 2016; Jiang & Lu, 2018), allowing agents to transmit observations, intentions, and advice to facilitate decision-making and coordination (Zhu et al., 2022; Qu et al., 2021). This approach, analogous to human language or verbal exchanges (Havrylov & Titov, 2017; Baker et al., 1999), has been widely employed in MARL to enhance collaboration. However, dependence on direct channels introduces significant computational and memory overheads (Roth et al., 2006), which makes it challenging to implement in certain scenarios. Examples are tasks without communication channels or decentralized frameworks (Oliehoek et al., 2008; Kraemer & Banerjee, 2016).
@@
To address these challenges associated with ToM methods, we introduce Implicit Channel Protocol (ICP), a novel framework that allows agents to communicate through communication protocols in implicit communication likewise how it was done in explicit communication. ICP leverages a subset of actions, denoted as the scouting actions, which have no or uniform effects on environment dynamics. A centralized mapping between information and these scouting actions is established to encode and decode the messages. Agents exchange information by deliberately taking scouting actions, forming an implicit communication channel. We further demonstrate how agents’ strategies are trained on this channel, including training with a randomly initialized information map and training with a delayed information map.
We validate the effectiveness of ICP through comprehensive experiments on the tasks of Guessing Numbers, Revealing Goals, and Hanabi (Bard et al., 2020). These environments share a common characteristic: they lack direct communication but agents must collaboratively make decisions to achieve shared rewards. This setting introduces significant challenges, including sparse and delayed reward feedback, along with difficulty in credit assignment both temporally and among agents. Despite these hurdles, our experiments on Guessing Numbers and Revealing Goals demonstrate that ICP significantly enhances performance, through more efficient information transmission, compared to baseline methods. In Hanabi, which is a popular card game played by humans, our approach achieved an average score of $24.91$ out of $25$, which surpasses the best available learning algorithm which obtains $23.81$.
@@ 2 Background and Related Work
The cooperative multi-agent problem can be formulated as a Dec-POMDP game (Bernstein et al., 2002), which is described by the tuple $G=(S,U,N,\mathcal{T},O,R,\gamma)$, where $S$ represents the true global state of the environment. At each discrete time step $t$, every agent $i\in N:=\{1,\ldots,n\}$ selects an action $u_{i}$ from its action space $U_{i}$. The state transition function $\mathcal{T}(s^{\prime}|s,u):S\times U\times S\rightarrow P(S)$ governs the transition of states, where $u=(u_{1},\ldots,u_{n})$ denotes the joint action. In POMDP, the global state remains inaccessible, and each agent $i$ can only perceive its individual observation $o_{i}$ through the observation function $O_{i}(s):S\times N\rightarrow O$. $R_{i}(s,u_{i}):S\times U_{i}\rightarrow R$ represents the reward function for each agent $i$. In the cooperative scenarios, all agents share a common reward function $R(s,u):S\times U\rightarrow R$, known as team reward. The objective for each agent is to maximize the expected return, which makes effective cooperation among all agents necessary.
@@
Communication among agents is critical for effective collaboration in MARL. Most researches in this field focused on explicit communication protocols (Tucker et al., 2022; Peng et al., 2017; Kong et al., 2017; Pesce & Montana, 2020; Kim et al., 2019; Wang et al., 2020; Freed et al., 2020b; a; Gupta et al., 2023; Foerster et al., 2016), where agents exchange messages containing critical information. Additionally, techniques like attention mechanisms or graph-based methods that used to build more effective communication connection (Jiang & Lu, 2018; Das et al., 2019; Jiang et al., 2018; Sukhbaatar et al., 2016; Chen et al., 2024; Tucker et al., 2022) and intention sharing or theory of mind (ToM) reasoning to transmit more useful information (Wang et al., 2021; Kim et al., 2020; Qu et al., 2021), further enhancing the effectiveness of MARL systems.
Explicit communication on direct channels often incurs significant communication overhead and may not be available in environments with limited bandwidth and high communication costs. Implicit communication, where agents convey information through their behaviors, has been explored as an alternative approach (Li et al., 2024; Tian et al., 2023; Li et al., 2021; Shaw et al., 2022; Grupen et al., 2022). ToM-based methods represent one main research strand for implicit communication, where agents infer the intentions and beliefs of others to anticipate actions and adjust strategies accordingly. Methods in Rabinowitz et al. (2018); Tian et al. (2020); Nguyen et al. (2020); Zhao et al. (2023) aim to enhance coordination by modeling and predicting others’ behavior based on observed actions. However, ToM-based approaches can be computationally intensive and may struggle to accurately infer intentions in complex environments.
@@ 3 Setting
In an idealized situation, scouting actions serve to influence the observation mapping function, which allows the agents to gather critical information from the environment to improve their decision-making and coordination. A scouting action might reveal information about an agent’s surroundings or the state of other agents without changing the environment unpredictably. For instance, in the game of Hanabi, Hint actions reveal information to other players while consuming an information token, thereby uniformly altering the game state. Similarly, in StarCraft II, scouting the map with Scan from Terrans requires depleting energy, which impacts the environment and provides new information in the agent’s observations.
@@ 4 Information Carried by Scouting Actions
Scouting actions carry two types of information.
@@
    Information reflected through the environment. When an agent performs a scouting action, it influences the observation function, and the new observations provide information. This type of information depends on the observation function, the states, and the joint actions of all agents.
@@
    Information reflected through the choice of the scouting action. The specific scouting action chosen by an agent can carry intentional meaning. This type of information depends only on other agents’ scouting action policy. If the receiver agent cannot understand the sender agent’s intention, no information will be transmitted.
General multi-agent reinforcement learning (MARL) methods tend to focus more on the first type of information. In Dec-POMDPs without modeling other agents’ strategies, both partial observation of states and other agents’ policies introduce uncertainty. In the exploration phase of online MARL training, the second type of information cannot be utilized due to the unknown policies of others. Therefore, MARL methods mostly learn to use the information reflected through the environment.
For the second type of information which is not limited by observation function and states, when agents’ intention is understood correctly, it is more fixable and could be more useful for decision-making and cooperation. Theory of mind (ToM) methods utilize the second type of information by modeling strategies or inferring the intention of other agents. By estimating the internal states of other agents, an agent can interpret the intentional choices behind their scouting actions. One approach is to build an explicit belief model that infers the state from observed actions to achieve communication. However, this method incurs a computational cost that grows exponentially with the size of the game and the number of agents. Moreover, the precise state distribution information is often difficult to fully utilize. Another approach involves training a neural network to map actions to states or intentions, which requires exponentially more data compared to action policy training. Since this method does not produce explicit inferences, the resulting information is inherently biased, which could negatively impact the agent’s decision-making process.
We propose a simpler and more effective solution to take advantage of the second type of information conveyed by scouting actions while also mitigating the computational overhead and inaccuracies of ToM methods. By establishing a shared communication protocol where all agents follow the same strategy for selecting scouting actions, the intended information can be easily decoded from the chosen action. This approach eliminates the need for state estimation and intention inference required in ToM methods. In our setting, this communication protocol has the following benefits.
@@
    Low Computational Overhead and Accurate Information Transmission. By having all agents follow the same strategy for selecting scouting actions, there is no need for additional computational inference or training of a separate inference model to deduce intentions. The inverse of the strategy can be used to directly interpret the information, resulting in low computational overhead and accurate transmission of intentions.
@@
    Efficient Communication Enabled by Broadcasted Information. Scouting actions are observable by all agents in the system, allowing any information conveyed through these actions to be inherently broadcasted. This characteristic facilitates the use of efficient embedding techniques, such as the hat mapping method, to achieve more effective communication.
@@
    Independence from Observation-Based Information. When different embeddings are used, the information conveyed by altering the observation function may differ, making the information carried by scouting actions not necessarily correlated with the environment-based information. Both types of information can be optimized concurrently to enhance cooperation and decision-making.
@@ 5 Strategy Training on Constructed Implicit Channel
To establish a shared communication protocol, we propose Implicit Channel Protocol (ICP), a new framework for agents’ implicit communication. In this framework, at each step, agents choose whether to send information. If an agent chooses not to send information, a regular action $u^{r}$ will be taken according to the action policy. If otherwise an agent wishes to convey message $m_{i}\in M$, it will output a scouting action $u^{s}=\mathcal{P}(m_{i})$, which maps $m_{i}$ through the centralized mapping mechanism $\mathcal{P}$. This mechanism $\mathcal{P}$ needs to be decodable, which provides output messages from scouting actions. By allowing agents to send information through encoded scouting actions and other agents to decode the messages from the globally observed scouting actions, an implicit communication channel is constructed.
@@
Equation (1) highlights three essential components of the framework: the agents’ action policy $\pi$, the message strategy $\phi$, and the centralized mapping mechanism $\mathcal{P}$. To develop these components, we investigate two different methodologies. The first one is training with a random initial information map and the second one is training with a delayed information map. Both approaches construct an implicit communication channel within the environment, with which the agents exchange information and achieve implicit coordination. In the rest of the section, we will discuss the details of these approaches.
@@ 5.1 Strategy Training with Random Initial Map
In this approach, we first establish a one-to-one mapping between information and actions using a randomly initialized embedding, that makes the size of the message the same as the size of the scouting action space $M=\{1,...,|U^{s}|\}$. Once this one-to-one mapping is set, the inverse of $\mathcal{P}$ will obtain the original message $m_{i}=\mathcal{P}^{-1}(u_{i}^{s})$.
In our implementation, we use a value-based method for training the action policy, incorporating parameter sharing and value decomposition techniques to facilitate effective coordination among agents. For the information strategy, we use the communication gradient method (Foerster et al., 2016), which is designed to optimize communication protocols within a predefined channel. Both the action policy and the information strategy are parameterized as Q-networks, for action policy $\pi_{i}(\tau_{t,i})=\text{argmax }Q^{\theta_{1,i}}(\tau_{t,i},u_{i})$, and message strategy $\phi_{i}(\tau_{t,i})=\text{argmax }Q^{\theta_{2,i}}(\tau_{t,i},m_{i})$. Given the value function $V^{\pi_{i},\phi_{i},\mathcal{P}}(s)$, the Q-functions are defined by $Q^{\pi_{i}}(\tau_{t,i},u_{i}):=\mathbb{E}_{m\sim\phi,\mathcal{P}}[r_{t}(s_{t},u_{t},s_{t+1})+\gamma V^{\pi_{i},\phi_{i},\mathcal{P}}],$ and $Q^{\phi_{i}}(\tau_{t,i},m_{i}):=\mathbb{E}_{u^{r}\sim\pi,\mathcal{P}}[r_{t}(s_{t},u_{t},s_{t+1})+\gamma V^{\pi_{i},\phi_{i},\mathcal{P}}],$ where $\pi=(\pi_{1},...,\pi_{n})$, $\phi=(\phi_{1},...,\phi_{n})$ and $m=(m_{1},...,m_{n})$.
@@
This method ensures a continuous relaxation between $\phi$ and $\vec{m}$. As a result of this relaxation, the game becomes fully differentiable and can be trained using the backpropagation algorithm. By enabling gradients to flow smoothly through the communication process, this approach facilitates more effective and efficient training of agents. Comparing the communication gradient implementation in DIAL (Foerster et al., 2016), we adopt this Gumbel-Softmax technique to replace the Discretize/Regularize Unit (DRU). This substitution allows agents to send discrete and non-binary messages during the learning phase, thereby enhancing the richness and accuracy of the information exchanged.
@@ 5.2 Strategy Training with Delayed Map
In the training process with a delayed map, we begin by learning the information strategies and action policies using a direct communication channel. This initial phase provides agents with a flexible environment that allows for direct communication and enable them to develop their strategies more effectively. Once these strategies are established, we determine the mapping mechanism that maps the messages to actions. Finally, we fine-tune both the action policies and information strategies in the ICP framework to ensure optimal performance within the constraints of the final environment.
Given that the information strategies were initially trained using a direct communication channel, there may be discrepancies between this channel and the implicit communication channel constructed by scouting actions, particularly in terms of capacity and whether the channel is discrete or noiseless. To minimize any potential performance loss resulting from these differences, we can limit the capacity of the direct channel or apply efficient embedding techniques (such as the hat mapping method) during the message-to-action mapping process. These steps help to align information strategies developed in the direct channel with the constraints of the implicit channel and ensure messages are decodable.
During implementation, we utilize standard explicit communication algorithms designed for direct channels, such as RGMComm, to pre-train the information strategy. Then we use two kinds of mapping methods, including simple one-to-one mapping and the hat mapping method (Brown & Tanton, 2009; Bushi, 2012; Butler et al., 2009; Feige, 2004; Havil, 2011; Winkler, 2002) to obtain new information strategies. The one-to-one mapping ensures that each message corresponds directly to an action, while the hat mapping allows an agent to communicate efficiently with multiple receivers. Finally, we fine-tune both the information strategy and the action policy to optimize performance within the implicit communication framework.
@@
In the pre-training phase, we utilize a direct communication channel to facilitate the learning of the communication strategies. During this phase, the message size can be larger than the scouting action space, allowing for more complex and detailed communication between agents. This flexibility enables the agents to explore a wider range of communication strategies and coordination behaviors, which are crucial for solving complex tasks in partially observable environments. However, after pre-training, the message strategy needs to transition to the implicit communication framework and it is necessary to compress these messages without loss of information. For one-to-one mapping, this mapping requires that the message size matches the scouting action space exactly. It ensures a direct and consistent mapping between the learned messages and the corresponding actions.
@@
When the local observations of different agents’ have large overlaps, the hat principle-based mapping method utilizes broadcast channels to achieve a single transmission of messages to multiple receivers. The method is inspired by the multi-color hat guessing game, where players use logical deduction based on limited communication to maximize team success. In this game, players can observe the hat colors of others but not their own, and they must guess their own hat color using a pre-defined strategy. A well-known strategy involves each player leveraging the sum of the hat colors they observe, modulo the total number of possible colors. For example, in an 8-color version of the game, each player can compute the sum of the colors they observe (assigned numerical values) and use a pre-determined rule, such as calculating the sum modulo 8, to make a logical deduction about their own hat color.